# Bloch oscillation with a diatomic tight-binding model on quantum computers

### v1

@@ II Two-band tight-binding model
We consider a simple diatomic tight-binding model on a one-dimensional, open-ended chain, as shown in Fig. 1. Two types of atoms are placed in each unit cell with type-A on even sites and type-B on odd sites. An electron can hop between type-A and type-B atoms within a single cell with hopping constant $\triangle_{a}/4$, or between two neighboring cells with hopping constant $\triangle_{b}/4$. In the presence of a time-dependent linear electric potential (which corresponds to a spatially-uniform electric field along the chain), the second quantization representation of tight-binding model Hamiltonian has three contributions,
@@ II.1.1 Jordan-Wigner representation
[…]
@@ IV Numerics
In this section, we show some numerical attempts on freely available IBM hardware (ibm_brisbane and ibm_sherbrooke) through cloud quantum computing access. In the following, we show the simulation results compared with exact solutions for non-interacting single-particle state cases with two and three qubits.
@@
Figure 14: Demo plots of probability distribution $|\psi(2,t)|^{2}$ on IBM hardware (black dots) without error mitigation vs. exact numerical result (red solid) for a 2-qubit and $4$-site finite system under the parameters $\Delta_{a}=5$ , $\Delta_{b}=1$, $F=1.5$. The time evolution step is $\delta t=0.02$.
Figure 15: Similar to Fig. 14, but for a 3-qubit and $8$-site finite system.
@@ V Summary and outlook
The two-band diatomic tight-binding 1D model, which is relatively simple but still rich in physics content, is used to demonstrate the effectiveness of quantum simulation under statevector basis representation. Exact expressions for quantum circuits of unitary time evolution of hopping terms, electric field term, and interaction term are obtained. The few-body dynamics of the model can be simulated with only a few qubits and simple quantum circuits. Numerical tests on IBM hardware show excellent agreement on a 2-qubit and 4-site system, even without error mitigation. On the other hand, testing results on the same IBM hardware are not satisfactory when the number of qubits is greater than two. Unfortunately, without significant effort on error mitigation and error correction, quantum simulation on the current noisy intermediate-scale quantum (NISQ) hardware are too noisy with more qubits.
The extension to higher dimensions in the two-band tight-binding model is fairly straightforward, by adding extra sets of quantum registers for extra dimensions. This is outlined in Appendix B.
@@ Appendix A Basic properties of diatomic non-interacting tight-binding model
$\phi^{(\nu)}_{k}(l)$ =$=\begin{cases}\frac{1}{\sqrt{2}}D^{(\nu)}(k)e^{-ikl},&l=2n,\\ \frac{1}{\sqrt{2}}e^{-ikl},&l=2n+1,\end{cases}$
[…]
@@
[…]
@@
Figure 17: Expectation value of electron position in a $20$-site finite system with hopping parameters $\Delta_{a}=5$, $\Delta_{b}=1$ and constant electric field $F=1$. Top: type-A (black) vs. type-B (red) as defined in Eq.(51). Bottom: both types as an average defined in Eq.(52).
Figure 18: Electron probabilities $|\Psi_{A}(t)|^{2}$ (black) vs. $|\Psi_{B}(t)|^{2}$ (red) as defined in Eq.(54) for a $20$-site finite system under the parameters: $\Delta_{a}=5$, $\Delta_{b}=1$, $F=1$.
Figure 19: Expectation value of electron momentum for type-A wave $\langle k_{A}(t)\rangle$ (black) vs. type-B wave $\langle k_{B}(t)\rangle$ (red) for a $20$-site finite system with parameters: $\Delta_{a}=5$, $\Delta_{b}=1$, $F=1$.
@@
They are plotted in Fig. 17 with a gaussian initial wavefunction,
@@
We see that the Bloch oscillation in two-band tight-binding system with unequal hopping exhibits a more complex pattern than that of equal coupling: not only the type-A and type-B waves oscillate separately moving against each other (almost completely out of phase), but there is a tunneling effect in the expectation value of site-$l$ of the entire 1D chain also oscillates, as indicated in the bottom panel of Fig. 17.

### v2

@@ II Two-band tight-binding model
We consider a simple diatomic tight-binding model on a one-dimensional, open-ended chain, as shown in Fig. 1. Two types of atoms are placed in each unit cell with type-A on even sites and type-B on odd sites. An electron can hop between type-A and type-B atoms within a single cell with hopping constant $\triangle_{a}/4$, or between two neighboring cells with hopping constant $\triangle_{b}/4$. The simplest realization of such an (AB)(AB)(AB)-(AB) chain-like system in nature with different hoppings between intra- and extra-cell would be the polyyne type of periodic $C-C\equiv C-C$ structure with alternating single and triple bonds in-between carton chains. In the presence of a time-dependent linear electric potential (which corresponds to a spatially-uniform electric field along the chain), the second quantization representation of tight-binding model Hamiltonian has three contributions,
@@ II.1.1 Jordan-Wigner representation
$|\Psi(t)\rangle=\mathcal{T}\left[e^{-i\int_{0}^{t}dt^{\prime}\hat{H}_{qs}(t^{\prime})}\right]|\Psi(0)\rangle=\sum_{l=0}^{N-1}\psi(l,t)c_{l}^{\dagger}|0\rangle,$ (9)
@@ IV Numerics
In this section, we show some numerical attempts on freely available IBM hardware (ibm_brisbane and ibm_sherbrooke) through cloud quantum computing access. In the following, we show the simulation results compared with exact solutions for non-interacting single-particle state cases with two and three qubits. We also show the result for two interacting particles on a 4-site chain with four qubits (two sets of two-qubit, one set for each particle).
@@
Figure 14: Demo plots of probability distribution $|\psi(2,t)|^{2}$ on IBM hardware (black dots) without error mitigation vs. exact numerical result (red solid) for a 2-qubit and 44-site finite system under the parameters $\Delta_{a}=5$ , $\Delta_{b}=1$, $F=1.5$. The time evolution step is $\delta t=0.02$.
Figure 15: Similar to Fig. 14, but for a 3-qubit and 88-site finite system.
With interaction turned on, 4-qubit quantum circuit of lowest order Trotterization approximation of Eq.(32) for two interacting particles on a 4-site periodic chain is shown in Fig. 16. The initial state is chosen as a spike at site $l_{1}=1$ and $l_{2}=2$ for particle-1 and particle-2 respectively,
$\psi(l_{1},l_{2},0)=\delta_{l_{1},1}\delta_{l_{2},2}.$ (38)
A comparison of the exact solution of $\psi(1,2,t)$ and the results from running on IBM hardware is shown in Fig. 17 with four qubits. Similar to the 3-qubit non-interacting single particle result shown in Fig. 15, the 4-qubit two-particle interacting system is also way off the exact result even with error mitigation effort. To highlight the effects of the interaction, the exact solution of $\psi(1,2,t)$ with interaction turned off ($V=0$) is also plotted in Fig. 17. At the limit of $V=0$, the two-particle wave function is simply the product of two single particle wave function:
$\psi(l_{1},l_{2},t)\stackrel{{\scriptstyle V\rightarrow 0}}{{\rightarrow}}\psi(l_{1},t)\psi(l_{2},t).$ (39)
Figure 16: Quantum circuit of the lowest order $U(\delta t)$ in Eq.(32) for 4-qubit (two sets of 2-qubit) two interacting particles on 4-site periodic chain.
Figure 17: Demo plots of $|\psi(1,2,t)|$ on IBM hardware (black dots) without error mitigation, results with error mitigation (purple dots) of Zero Noise Extrapolation method [28] vs. exact numerical result (red solid) for a 4-qubit (two sets 2-qubit) and 44-site two interacting particles system under the parameters $V=10$, $\Delta_{a}=5$ , $\Delta_{b}=1$, $F=1.5$. The time evolution step is $\delta t=0.02$. The $|\psi(1,2,t)|$ with interaction turned off ($V=0$, blue dashed) is also plotted as a comparison.
@@ V Summary and outlook
The two-band diatomic tight-binding 1D model, which is relatively simple but still rich in physics content, is used to demonstrate the effectiveness of quantum simulation under statevector basis representation. Exact expressions for quantum circuits of unitary time evolution of hopping terms, electric field term, and interaction term are obtained. The few-body dynamics of the model can be simulated with only a few qubits and simple quantum circuits. Numerical tests on IBM hardware show excellent agreement on a 2-qubit and 4-site system, even without error mitigation. On the other hand, testing results on the same IBM hardware are not satisfactory when the number of qubits is greater than two. Unfortunately, without significant effort on error mitigation and error correction, quantum simulation on the current noisy intermediate-scale quantum (NISQ) hardware are too noisy with more qubits. For the interacting case, even for the two particles on 4-site with two sets of 2-qubit, the testing result on IBM hardware is not satisfactory. The reason of failure is not completely clear at this point. We suspect that the deep quantum circuit may be the primary cause. Using non-interacting 3-qubit case as an example, in terms of $U1$, $U3$ and $CX$ gates, the depth of decomposed quantum circuit $U(\delta t)$ in Fig. 13 is 25 layers, with 4 $U1$ gates, 13 $U3$ gates and 14 expensive $CX$ gates. Consequently, $U(10\delta t)$ would involve 250 layers which may become too challenging for current quantum hardware with decoherence time in the range of hundred microseconds without significant efforts on error mitigation and error correction.
The extension to higher dimensions in the two-band tight-binding model is fairly straightforward, by adding extra sets of quantum registers for extra dimensions. This is outlined in Appendix B. The extension to include multiple orbits per site could possibly be dealt with in a similar way to that for higher dimensions.
@@ Appendix A Basic properties of diatomic non-interacting tight-binding model
$\phi^{(\nu)}_{k}(l)$ $=\begin{cases}\frac{1}{\sqrt{2}}D^{(\nu)}(k)e^{-ikl},&l=2n,\\ \frac{1}{\sqrt{2}}e^{-ikl},&l=2n+1,\end{cases}$
$\text{ with }D^{(\nu)}(k)=-\frac{4\epsilon_{\nu}(k)}{\Delta_{a}e^{ik}+\Delta_{b}e^{-ik}}.$ (47)
@@
$E^{(\nu)}_{\alpha}\simeq 2F\alpha+\frac{1}{\pi}\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}dp\left[\epsilon_{\nu}(p)-FX_{\nu\nu}(p)\right],\ \ \alpha\in\mathbb{Z}.$ (50)
@@
Figure 19: Expectation value of electron position in a $20$-site finite system with hopping parameters $\Delta_{a}=5$, $\Delta_{b}=1$ and constant electric field $F=0.2$. Top: type-A (black) vs. type-B (red) as defined in Eq.(53). Bottom: both types as an average defined in Eq.(54).
Figure 20: Similar to Fig. 19, but with $F=1$.
Figure 21: Electron probabilities $|\Psi_{A}(t)|^{2}$ (black) vs. $|\Psi_{B}(t)|^{2}$ (red) as defined in Eq.(56) for a $20$-site finite system under the parameters: $\Delta_{a}=5$, $\Delta_{b}=1$, $F=0.2$ (upper panel) and $F=1$ (lower panel).
Figure 22: Expectation value of electron momentum for type-A wave $\langle k_{A}(t)\rangle$ (black) vs. type-B wave $\langle k_{B}(t)\rangle$ (red) for a $20$-site finite system with parameters: $\Delta_{a}=5$, $\Delta_{b}=1$, $F=0.2$ (upper panel) and $F=1$ (lower panel).
@@
They are plotted in Fig. 19 and Fig. 20 with a gaussian initial wavefunction,
@@
We see that the Bloch oscillation in two-band tight-binding system with unequal hopping exhibits a more complex pattern than that of equal coupling: not only the type-A and type-B waves oscillate separately moving against each other (almost completely out of phase), but there is a tunneling effect in the expectation value of site-$l$ of the entire 1D chain that also oscillates, as indicated in the bottom panel of Fig. 19 and Fig. 20.